OpenStudy (anonymous):

Help pleaseeeeee!!! Explain the grouping method of factoring. Describe a scenario when the grouping method would be preferred over other methods and provide an example of this type of problem.

OpenStudy (anonymous):

a^2 b + c^2 d - abc - acd a^2 b - abc + c^2 d - acd ab(a - c) + cd(a - c) (ab + cd)(a - c) Hope this helps.

OpenStudy (anonymous):

I've already read that on a different question but I still don't get it. What does that mean? Is there supposed to be a symbol between "a^2" and "b" and "c^2" and "d"??? I'm terrible at Algebra and everything goes over my head

OpenStudy (anonymous):

i know thats why i looked it up. cause im in 8th grade and really bad at math. well kinda i can do math at about a 9th grade level.

Directrix (directrix):

@francyhive Here's an example of using factoring by grouping to factor a trinomial. "Bust the b" technique for factoring 4 x² -1x + 18 a = 4 and c = 18 and b = -1. The task on "bust the b" is to find numbers that mutiply to a*c AND add to b. In this case, numbers that multiply to 4*18 and add to -1. 14 * 18 breaks aparts into the product of prime factors : 2*2*2*3*3 Look at those five factor of 14 * 18 and try to put them in the form of 2 numbers that differ by -1. 2*2*2 and 3*3 appear to differ by 1. That would be 8 and 9. "b" has been "busted into 8 and - 9 whose sum is -1 which equals "b" in this problem. 4x - x+ 18 = 4 x² + 8 x - 9 x - 1. The "b" has been busted. Factoring by grouping comes next. 4 x² + 8 x - 9 x - 18 = 4x ( x + 2) - 9 ( x + 2 ) = --> On this step, (x + 2) is the common factor of the expression. [ (x + 2) ] ( 4x - 9 ) = ( x + 2 ) ( 4x - 9)